# Flipping Out: Role of Arginine in Hydrophobic Polymer Collapse

### v2

@@ Abstract
Arginine has been a mainstay in biological formulation development for decades. To date, the way arginine modulates protein stability has been widely studied and debated. Here, we employed a hydrophobic polymer to decouple hydrophobic effects from other interactions relevant to protein folding. While existing hypotheses for the effects of arginine can generally be categorized as either direct or indirect, our results indicate that direct and indirect mechanisms of arginine co-exist and oppose each other. At low concentrations, arginine was observed to stabilize hydrophobic polymer collapse via a sidechain-dominated direct mechanism, while at high concentrations, arginine stabilized polymer collapse via a backbone-dominated indirect mechanism. These findings highlight the modular nature of the widely used additive arginine, with relevance in the design of stable biological formulations.
keywords:
Hydrophobic interactions, formulation design, excipient, osmolyte, protein stability, protein folding
[…]
{tocentry}
@@
Maintaining native protein structures in biological formulations poses a challenge, and is commonly addressed by strategic additive incorporation.1, 2, 3 Arginine stands out as a frequently employed additive in such formulations, spanning both therapeutic proteins4 and vaccines.5, 6 Arginine has been widely used as an aggregation suppressor, an agent for protein refolding, a cryoprotectant during lyophilization, and in protein purification. 7, 8, 9 Once hailed as a universal stabilizer, emerging studies paint a foggier picture of the effects of arginine. In some settings, the presence of arginine has accelerated the aggregation 10, 11, 12, denaturation13, 14, 15, and inactivation16 of certain proteins and viruses. Additional studies have found that arginine increases or decreases protein melting temperature depending on concentration 17, 18. Hence, the existing literature on arginine reveals a lack of a cohesive understanding of its multi-faceted effects on protein stability.
@@
The molecular mechanism for arginine effects on protein stability remains elusive. In general, additive effects on protein stability are thought to be either direct or indirect. Direct mechanisms involve direct protein-additive interactions that drive stabilization, while indirect mechanisms influence protein stability by modulating the surrounding solvent structure. It remains debated whether arginine acts primarily via a direct or indirect mechanism. Several studies called attention to direct interactions between arginine and aromatic residues 7, 19, 20, acidic residues 12, 17, and hydrophobic moieties 21, 22. Other studies have proposed clusters of free arginine molecules in solution enable the crowding out of protein-protein interactions 23, 20, 24, 17, or alteration of hydration shell water dynamics.25, 26 The wide range of observations related to the role of arginine on protein stability suggests that arginine harbors diverse, context-dependent mechanisms.
@@
To elucidate the mechanisms through which arginine influences protein stability, this study focuses on its effects on hydrophobic interactions. Hydrophobic interactions are key in several biologically relevant phenomena, including protein folding and stability.27, 28, 29, 30, 31, 32, 33, 34, 35 Additives in solutions have been observed to modulate the strength of hydrophobic interactions, and in turn, the stability of proteins.36, 37, 38, 39, 40, 41, 42 For example, simulation studies have shown that trimethylamine N-oxide (TMAO) has a negligible effect on or strengthens hydrophobic interactions,38, 43, 44, 45, 46, 47 while hydrophobic interactions are weakened in urea solutions.48, 49, 37, 50, 51, 39, 40, 52 Indeed, the effects of urea and TMAO on hydrophobic interactions are consistent with their experimentally-observed roles as a protein denaturant and stabilizer, respectively.53, 54, 55, 41
@@
Molecular dynamics (MD) simulations have provided valuable insights towards understanding these mechanisms as they relate to hydrophobicity. Several studies have highlighted the utilization of a hydrophobic polymer model for describing the role of solvent and additives on protein-like collapse. 56, 40, 57, 58, 39, 59 The use of a hydrophobic polymer model enables the decoupling of additive effects on hydrophobic vs other interactions, which is challenging in experiments. In the present study, we utilize MD simulations of a hydrophobic polymer to characterize the effects of arginine on many-body hydrophobic interactions pertinent to protein stability.
@@
We simulated a hydrophobic polymer in arginine solutions at different concentrations (Table S2). Replica exchange umbrella sampling60 (REUS) simulations were performed using GROMACS 2021.461, 62 with the PLUMED 2.8.0 63, 64 patch applied. The hydrophobic polymer was modeled as a linear coarse-grained chain with 26 monomers, where each monomer represents a $\mathrm{CH_{2}}$ unit with Lennard-Jones parameters $\sigma=0.373$ $\mathrm{nm}$ and $\epsilon=0.5856$ $\mathrm{kJ/mol}$38. Chloride (Cl-) counterions were added to achieve charge neutrality. The TIP4P/200565 model was used to describe water, and the CHARM22 force field was used for arginine and Cl- 66. Lorentz-Berthelot mixing rules67 were used to calculate non-bonded interactions between different atom types, except for polymer-water oxygen interactions (see SI for details).
@@
The potential of mean force (PMF) along the radius of gyration of the polymer was calculated as ${W(R_{g})=-k_{B}Tln(P(R_{g}))}$. Biased probability distributions were reweighted according to the Weighted Histogram Analysis Method (WHAM).68 The free energy of polymer unfolding ($\mathrm{\Delta G_{u}}$) was calculated according to:
@@
Fig. 1a shows the PMF along the $R_{g}$ reaction coordinate. In all solutions, free energy minima were observed at approximately 0.4 and 0.8 nm (labeled as III in Fig. 1g), along with a prominent free energy barrier at $\sim$0.6 nm representing the transition between folded and unfolded states. (Fig. 1a). In pure water, hydrophobic collapse is unfavorable, with the unfolded state favored by $\sim$0.3 kT. In contrast, at all arginine concentrations, the folded state of the polymer is favored relative to pure water, and a monotonic increase in $\Delta G_{u}$ is observed (Fig. S4). An additional barrier at $\sim$0.45 nm was identified separating two folded states, labeled as I and II in Fig. 1a.
The folded polymer ensemble in arginine solutions exhibits free energy minima corresponding to globular ($\sim$0.4 nm) and hairpin-like ($\sim$0.5 nm) configurations (Fig. 1g). Arginine clusters encapsulating the hydrophobic polymer are observed in each state (Fig. S6). We propose that the free energy barrier separating these two states arises from an energetic penalty associated with breaking these encapsulating clusters. Such a mechanism is similar to that observed by Li et al.22, who observed arginine-mediated suppression of hydrophobic association.
The PMF was decomposed into individual components to further investigate the role of arginine in polymer collapse. Following the methods outlined by several others,69, 70, 42, 71 the PMF was decomposed as
$\begin{split}W(R_{g})&=W_{vac}(R_{g})+W_{cav}(R_{g})\\ &\quad+E_{pw}(R_{g})+E_{pa}(R_{g})\\ &\quad+E_{pc}(R_{g})\end{split}$ (2)
@@
Figure 1: PMF decomposition in 0.0 M, 0.25 M, 0.5 M, and 1.0 M arginine solutions. (a) The PMF obtained along the $R_{g}$ reaction coordinate, $W(R_{g})$, (b) vacuum contribution, $W_{vac}$, (c) cavitation contribution, $W_{cav}$, (d) polymer-water interactions, $E_{pw}$, (e) polymer-arginine interactions, $E_{pa}$, and (f) polymer-chloride interactions, $E_{pc}$. (g) Representative configurations along the reaction coordinate as denoted in (a) as I, II, and III. (h) Changes in overall free energy of unfolding ($\Delta\Delta G_{u}$), cavitation contribution ($\Delta\Delta G_{cav}$), polymer-water interactions ($\Delta\Delta E_{pw}$), and polymer-arginine interactions ($\Delta\Delta E_{pa}$). Increasing arginine concentration is denoted by increased shading (light to dark) and is indicated by arrows in (a-f). The polymer in water alone is shown in black, where appropriate. Mean values were estimated from three replicate REUS simulations. Error bars are reported as described in the SI. All plots are normalized to 0 at $R_{g}=0.4$ nm, where appropriate.
@@
Fig. 1b-f shows the decomposition of the PMF into these various components. The vacuum component favors the folded state of the polymer, associated with favorable intrapolymer interactions and configurational entropy upon collapse (Fig. 1b). Because this component does not depend on the presence of arginine, a balance of the remaining components dictates the effect of arginine on hydrophobic polymer collapse.
@@
Near large idealized solutes or hydrophobic interfaces, water dewets and forms a vapor-liquid-like interface72, 73. Similar behavior may be expected for our hydrophobic polymer; hence, we computed $W_{cav}$, which depends on both the size and shape of the polymer and is related to the air-water surface tension70, 74, 69. The cavitation component favors the folded state (Fig. 1c), reflecting a strong hydrophobic driving force for polymer collapse56.
@@
Attractive polymer-water interactions become more favorable with increasing $R_{g}$ (Fig. 1d), indicating polymer-water attractive interactions oppose polymer collapse. It is worth noting that the free energy minima at $\sim$0.8 nm in the unfolded ensemble is observed in an aqueous environment (Fig. 1a), but not in vacuum (Fig. 1b). This minima arises due to favorable polymer-water interactions, consistent with prior MD simulations that observed water-mediated interactions drive large hydrophobic solutes apart.75, 76, 57. Additionally, sufficient dewetting of the hydrophobic polymer is a known bottleneck to collapse,77, 56, 69 resulting in the free energy barrier at $\sim$0.6 nm separating the folded and unfolded states.
@@
Attractive polymer-arginine interactions approach an energetic minima at $\sim$0.5 nm (Fig. 1e), giving rise to the global minimum observed in the overall PMF. Polymer-chloride interactions were observed on the order of thermal fluctuations ($<0.5$ kT), consistent with previous observations that chloride ions are depleted from the local domain of hydrophobic solutes.78, 37
@@
Fig. 1h shows the change in each component upon unfolding in arginine solution relative to that observed in water. In Fig. 1h, the first $\Delta$ arises from the difference between folded and unfolded states (e.g., $\Delta E=\langle E_{u}\rangle-\langle E_{f}\rangle$), and the second $\Delta$ arises from the free energy difference between arginine solution ($\Delta E_{arg}$) and water ($\Delta E_{wat}$) (e.g., $\Delta\Delta E=\Delta E_{arg}-\Delta E_{wat}$). With increasing arginine concentration, we observed an increasingly favorable cavitation component (Fig. 1h). Lin and Timasheff 79 previously connected protein stability with cavity formation and vapor-liquid surface tension. They proposed that the expansion of a protein-containing cavity requires more energy in the presence of an additive that increases surface tension at the cavity-solution interface. Experimentally, increasing arginine concentration has been observed to increase the vapor-liquid surface tension of aqueous solutions80, which may explain the dependence of $\Delta\Delta G_{cav}$ observed in the present study.
@@
The polymer-arginine contribution is observed to favor the folded polymer state at 0.25 M and 0.5 M arginine concentrations relative to in pure water (Fig. 1h). However, at 1.0 M arginine concentrations, polymer-arginine interactions promote polymer unfolding. The fact that high concentrations of arginine favored polymer unfolding is evidence that the direct mechanism of arginine-induced stabilization cannot fully explain the effect of this additive on hydrophobic polymer collapse.
To investigate potential competing effects of direct and indirect mechanisms, we combined the components of our PMF decomposition, delineating between those linked to direct effects (polymer-arginine and polymer-chloride; $\Delta\Delta G_{dir}$) and indirect effects (cavitation and polymer-water; $\Delta\Delta G_{ind}$) of arginine (Fig. 2a).

### v3

# Flipping Out: Role of Arginine in Hydrophobic Interactions and Biological Formulation Design

@@ Abstract
Arginine has been a mainstay in biological formulation development for decades. To date, the way arginine modulates protein stability has been widely studied and debated. Here, we employed a hydrophobic polymer to decouple hydrophobic effects from other interactions relevant to protein folding. While existing hypotheses for the effects of arginine can generally be categorized as either direct or indirect, our results indicate that direct and indirect mechanisms of arginine co-exist and oppose each other. At low concentrations, arginine was observed to stabilize hydrophobic polymer collapse via a sidechain-dominated direct mechanism, while at high concentrations, arginine stabilized polymer collapse via a backbone-dominated indirect mechanism. When adding partial charges to sites on the polymer, arginine destabilized polymer collapse. Further, we found arginine-induced destabilization of a model virus similar to direct-mechanism destabilization of the charged polymer, and concentration-dependent stabilization of a model protein similar to the indirect mechanism of hydrophobic polymer stabilization. These findings highlight the modular nature of the widely used additive arginine, with relevance in the design of stable biological formulations.
MTU] Department of Chemical Engineering, Michigan Technological University, Houghton, MI 49931, USA UMA] Department of Chemical Engineering, University of Massachusetts Amherst, MA 01003, USA MTU] Department of Chemical Engineering, Michigan Technological University, Houghton, MI 49931, USA UMA] Department of Chemical Engineering, University of Massachusetts Amherst, MA 01003, USA
1 Introduction
@@
Maintaining native protein structures in biological formulations poses a challenge, and is commonly addressed by strategic additive incorporation.1, 2, 3 Arginine stands out as a frequently employed additive in such formulations, spanning both therapeutic proteins4 and vaccines.5, 6 Arginine has been widely used as an aggregation suppressor, an agent for protein refolding, a cryoprotectant during lyophilization, and in protein purification.7, 8, 9 Once hailed as a universal stabilizer, emerging studies paint a foggier picture of the effects of arginine. In some settings, the presence of arginine has accelerated the aggregation,10, 11, 12 denaturation,13, 14, 15 and inactivation16, 17 of certain proteins and viruses. Additional studies have found that arginine mechanisms are dependent on concentration.18, 19, 20, 21 Hence, the existing literature on arginine reveals a lack of a cohesive understanding of its multi-faceted effects on protein stability.
@@
In general, additive effects on protein stability are thought to be either direct or indirect. Direct mechanisms involve direct protein-additive interactions, while indirect mechanisms influence protein stability by modulating the surrounding solvent structure. It remains debated whether arginine acts primarily via a direct or indirect mechanism. Several studies called attention to direct interactions between arginine and aromatic residues,7, 22, 23 acidic residues,12, 20, 24 and hydrophobic moieties.25, 26 Other studies have proposed clusters of free arginine molecules in solution enable the crowding out of protein-protein interactions,27, 23, 28, 20 or alteration of hydration shell water dynamics.29, 30 The wide range of observations related to the role of arginine on protein stability suggests that arginine harbors diverse, context-dependent mechanisms.
@@
To elucidate the mechanisms through which arginine influences protein stability, this study focuses on its effects on hydrophobic interactions. Hydrophobic interactions are key in several biologically relevant phenomena, including protein folding and stability.31, 32, 33, 34, 35, 36, 37, 38, 39 Additives in solutions are known to modulate the strength of hydrophobic interactions, and in turn, the stability of proteins.40, 41, 42, 43, 44, 45, 46 For example, simulation studies have shown trimethylamine N-oxide (TMAO) has a negligible effect on or strengthens hydrophobic interactions,42, 47, 48, 49, 50, 51 while these interactions are weakened in urea solutions.52, 53, 41, 54, 55, 43, 44, 56 Indeed, the effects of TMAO and urea on hydrophobic interactions are consistent with their experimentally-observed roles as a protein stabilizer and denaturant, respectively.57, 58, 59, 45
@@
Molecular dynamics (MD) simulations have provided valuable insights towards understanding these mechanisms as they relate to hydrophobicity. Several studies have highlighted the utilization of a hydrophobic polymer model for describing the role of solvent and additives on protein-like collapse. 60, 44, 61, 62, 43, 63 The use of a hydrophobic polymer model enables the decoupling of additive effects on hydrophobic vs other interactions, which is challenging in experiments. Additionally, comparison of arginine mechanisms on purely hydrophobic interactions allows indirect insights into experimentally-observed effects. For example, studies in which effects mirror what we observe from the hydrophobic polymer are likely to reflect a dominant mediation of hydrophobic interactions. In the present study, we utilize MD simulations of a hydrophobic polymer to characterize the effects of arginine on many-body hydrophobic interactions pertinent to protein stability, and contextualize these findings within larger-scale models for biological formulations.
Overall, we found arginine stabilizes hydrophobic polymer collapse at all concentrations under study. Interestingly, we discovered arginine sits on the edge of a mechanistic flip, balancing between direct- and indirect-dominated effects. As a consequence of this balance, we found subtle modulation of the polymer chemistry (via partial charge incorporation) changes arginine from a stabilizing additive to a destabilizer of polymer collapse. Consequently, in practical examples of formulation design, we observed arginine has variable effects on a model virus and protein.
2 Methods
2.1 Hydrophobic Polymer System Setup and Molecular Dynamics Simulations
@@
We simulated a hydrophobic polymer in arginine solutions at different concentrations (Table S2). All simulations were performed using GROMACS 2021.464, 65 with the PLUMED 2.8.0 66, 67 patch applied. The hydrophobic polymer was modeled as a linear coarse-grained chain with 26 monomers, where each monomer represents a $\mathrm{CH_{2}}$ unit with Lennard-Jones parameters $\sigma=0.373$ $\mathrm{nm}$ and $\epsilon=0.5856$ $\mathrm{kJ/mol}$42 Box dimensions were defined such that 1.5 nm of space separated the fully elongated polymer from the nearest box edge. For simulations with no polymer, the same box dimensions were used. Arginine was modeled in accordance with a pH of 7, resulting in a protonated sidechain. An equal number of arginine molecules and Cl- atoms were added to the box until the desired concentration was reached. The TIP4P/200568 model was used to describe water, and the CHARM22 force field was used for arginine and Cl-.69
All simulations were initially subject to energy minimization using the steepest descent algorithm. NVT equilibration was carried out for 1 ns at 300 K, followed by a 1 ns NPT equilibration at 300 K and 1 atm. During equilibration, temperature was controlled according to the V-rescale thermostat,70 while pressure was controlled via the Berendsen barostat.71 Following equilibration, NPT production runs were completed using the Nosé-Hoover thermostat72 and Parrinello-Rahman barostat.73 Production runs were completed for 20 ns for arginine/water systems, and between 50-250 ns per window for arginine/polymer/water replica exchange umbrella sampling (REUS) runs (Table 2). In all simulations, the Particle Mesh Ewald (PME) algorithm was used for electrostatic interactions with a cut-off of 1 $\mathrm{nm}$. A reciprocal grid of 42 x 42 x 42 cells was used with $\mathrm{4^{th}}$ order B-spline interpolation. A single cut-off of 1 $\mathrm{nm}$ was used for van der Waals interactions. The neighbor search was performed every 10 steps. Lorentz-Berthelot mixing rules74, 75 were used to calculate non-bonded interactions between different atom types, except for polymer-water oxygen interactions (see SI for details).
2.2 Replica Exchange Umbrella Sampling
REUS76 simulations were completed to sample the hydrophobic polymer conformational landscape in arginine solutions. The radius of gyration ($R_{g}$) of the hydrophobic polymer was used as the reaction coordinate to describe polymer folding/unfolding. 12 umbrella potential windows were centered between $R_{g}$ = 0.3 and $R_{g}$ = 0.9 nm, with a spacing of 0.05 nm. A force constant of $K$ = 5000 kJ/mol/nm2 was used in all windows, with the exception of the window centered at $R_{g}$ = 0.45, which used $K$ = 1000 kJ/mol/nm2 (see SI for details).
@@ 2.2 Replica Exchange Umbrella Sampling
The potential of mean force (PMF) along the radius of gyration of the polymer was calculated as ${W(R_{g})=-k_{B}Tln(P(R_{g}))}$. Biased probability distributions were reweighted according to the Weighted Histogram Analysis Method (WHAM).77 The free energy of polymer unfolding ($\mathrm{\Delta G_{u}}$) was calculated according to:
@@
We decomposed the PMF into individual components to further investigate the role of arginine in polymer collapse. Following the methods outlined by several others,78, 79, 46, 80 the PMF was decomposed as:
$W(R_{g})=W_{vac}(R_{g})+W_{cav}(R_{g})+E_{pw}(R_{g})+E_{pa}(R_{g})+E_{pc}(R_{g})$ (2)
@@
2.3 Preferential Interaction Coefficients
Distribution of arginine with respect to the polymer can be described via the preferential interaction coefficient ($\Gamma_{PA}$),81, 82, 83
where $\mu$ is the chemical potential, $m$ is the concentration and $W$, $P$, and $A$ refer to water, polymer, and an additive, respectively. This parameter is calculated in simulations using the two-domain formula84, 85, 86 given by:
2.4 Hydrogen Bond Analysis
Hydrogen bonds were calculated according to geometric criteria of a donor-acceptor distance of $r\leq$ 0.35 $\mathrm{nm}$ and donor-hydrogen-acceptor angle within 150° < $\theta$ < 180°.87 Hydrogen bond existence correlation functions for water-water and arginine-water interactions were estimated according to:88, 89, 90
where $h_{ij}\left(t_{0}\right)$ is equal to 1 if there is a hydrogen bond between groups $i$ and $j$ at time $t_{0}$, and 0 if no hydrogen bond is present. An average over all possible values of the time origin $t_{0}$ was taken over the last 5 ns production simulations.
2.5 Contact Coefficients
Contact coefficients ($CC$) give a measure of excipient preference for interacting with specific residues. Here, we computed $CC$s as described by Stumpe and Grubmüller91:
$CC_{X}=\frac{N_{X-A}}{N_{X-W}}\frac{M_{W}}{M_{A}}$ (6)
where $N_{X-A}$ and $N_{X-W}$ are the number of atomic contacts of protein residue $X$ with an additive or water molecules, respectively. Atoms were defined to be in contact if any pair of heavy atoms were within a 0.35 nm cut-off. $CC$s are normalized by the total number of additive atoms ($M_{A}$) and water atoms ($M_{W}$) in solution. Values of $CC_{X}~{}\textgreater~{}1$ indicate preferential interaction with an additive for residue $X$, while $CC_{X}~{}\textless~{}1$ denotes preferential interaction with water.
2.6 Protein and Virus Simulation Setup
Porcine parvovirus (PPV) and hen egg white lysozyme (HEWL) were used as large-scale models for formulation design. To model PPV, we constructed a surface model using 15 monomers from the published crystal structure (PDB: 1K3V).92 This was achieved by selecting three reference $C_{\alpha}$ atoms from the 5-fold pore of the viral surface, and aligning vector normal to these points with the z-axis of the simulation box. Box dimensions were 24 x 24 x 12 nm, with the capsid surface model extending approximately 9 nm in the z-direction. Arginine molecules were added to the exterior of the capsid surface to reach a target concentration of 0.25 M, and NaCl atoms were distributed throughout the box to both neutralize the system and reach a target concentration of 0.15 mM. To prevent diffusion of arginine molecules into the capsid interior, a wall was placed at z = 0 nm, and the capsid surface was positioned on top of this wall (Fig. S12). A position restraint with a force constant of 1000 kJ/mol/nm2 was applied to all capsid atoms within 1.5 nm from the wall. Capsid atoms between 1.5 nm and 3.0 nm from the wall were restrained with a force constant between 1-1000 kJ/mol/nm2, with atoms further from the wall scaled to a lesser extent. Capsid atoms greater than 3.0 nm from the wall were left fully flexible. PPV simulations were carried out in the NVT ensemble – first for 1 ns to reach the target temperature, followed 50 ns to allow the solvent to relax. 100 ns production runs were used for further analysis.
HEWL was modeled from an available crystal structure (PDB: 1LYZ),93 and simulated in solution with 0.1, 0.2, 1.0, and 2.0 M arginine. HEWL simulations were carried out in the NVT ensemble for 1 ns to equilibrate to the target temperature, followed by 1 ns NPT simulations to reach the target density. HEWL production runs were completed for 100 ns in the NPT ensemble. For both HEWL and PPV, the force field, thermostats, and barostats matched those used for the hydrophobic polymer systems.
2.7 Experimental Temperature Stability Studies
Liquid samples of PPV in arginine solutions were prepared in triplicates and were put either in a heat block at 60 °C 94 or in a fridge at 4 °C as the control samples. 72 hours later, the titer of PPV was determined by 3-(4,5-dimethylthiazol-2-yl)-2,5-diphenyltetrazolium bromide (MTT) colorimetric cell viability assay.95 Differential scanning fluorimetry (DSF) was performed to compute the hydrophobic exposure temperature (HET) of HEWL in arginine solutions.96, 97 More details for both experimental methods are included in the supporting information.
3 Results and Discussion
3.1 Arginine favors hydrophobic polymer collapse
Fig. 1a shows the PMF along the $R_{g}$ reaction coordinate. In all solutions, free energy minima were observed at approximately 0.4 and 0.8 nm (configurations labeled I and III in Fig. 1g), along with a prominent free energy barrier at $\sim$0.6 nm representing the transition between folded and unfolded states (Fig. 1a). In pure water, hydrophobic collapse is unfavorable, with the unfolded state favored by $\sim$0.3 kT. In contrast, at all arginine concentrations, the folded state of the polymer is favored relative to pure water, and a monotonic increase in $\Delta G_{u}$ is observed (Fig. S4). An additional barrier at $\sim$0.45 nm was identified separating two folded states, labeled as I and II in Fig. 1a.
The folded polymer ensemble in arginine solutions exhibits free energy minima corresponding to globular ($\sim$0.4 nm) and hairpin-like ($\sim$0.5 nm) configurations (labeled I and II in Fig. 1g). Arginine clusters encapsulating the hydrophobic polymer are observed in each state (Fig. S6). We propose that the free energy barrier separating these two states arises from an energetic penalty associated with breaking these encapsulating clusters. Such a mechanism is similar to that observed by Li et al.,26 who observed arginine-mediated suppression of hydrophobic association.
@@ 3.1 Arginine favors hydrophobic polymer collapse
Figure 1: PMF decomposition in 0.0 M, 0.25 M, 0.5 M, and 1.0 M arginine solutions. (a) The PMF obtained along the $R_{g}$ reaction coordinate, $W(R_{g})$, (b) vacuum component, $W_{vac}$, (c) cavitation component, $W_{cav}$, (d) polymer-water interactions, $E_{pw}$, (e) polymer-arginine interactions, $E_{pa}$, and (f) polymer-chloride interactions, $E_{pc}$. (g) Representative configurations along the reaction coordinate as denoted in (a) as I, II, and III. (h) Changes in overall free energy of unfolding ($\Delta\Delta G_{u}$), cavitation contribution ($\Delta\Delta G_{cav}$), polymer-water interactions ($\Delta\Delta E_{pw}$), and polymer-arginine interactions ($\Delta\Delta E_{pa}$). Increasing arginine concentration is denoted by increased shading (light to dark) and is indicated by arrows in (a-f). The polymer in water alone is shown in black, where appropriate. Mean values were estimated from three replicate REUS simulations. Error bars are reported as described in the SI. All plots are normalized to 0 at $R_{g}=0.4$ nm, where appropriate.
3.2 Decomposition of arginine effects on polymer folding
@@
Fig. 1b-f shows the decomposition of the PMF into various components. The vacuum component, $W_{vac}$, favors the folded state of the polymer, associated with favorable intrapolymer interactions and configurational entropy upon collapse (Fig. 1b). Because this component does not depend on the presence of arginine, a balance of the remaining components dictates the effect of arginine on hydrophobic polymer collapse.
@@
Near large idealized solutes or hydrophobic interfaces, water dewets and forms a vapor-liquid-like interface.98, 99 Similar behavior may be expected for our hydrophobic polymer; hence, we computed $W_{cav}$, which depends on both the size and shape of the polymer and is related to the vapor-liquid surface tension.79, 100, 78 The cavitation component favors the folded state (Fig. 1c), reflecting a strong hydrophobic driving force for polymer collapse.60
@@
Attractive polymer-water interactions become more favorable with increasing $R_{g}$ (Fig. 1d), indicating polymer-water interactions oppose polymer collapse. It is worth noting that the free energy minima at $\sim$0.8 nm in the unfolded ensemble is observed in an aqueous environment (Fig. 1a), but not in vacuum (Fig. 1b). This minima arises due to favorable polymer-water interactions, consistent with prior MD simulations that showed water-mediated interactions drive large hydrophobic solutes apart.101, 102, 61 Additionally, sufficient dewetting of the hydrophobic polymer is a known bottleneck to collapse,103, 60, 78 resulting in the free energy barrier at $\sim$0.6 nm separating the folded and unfolded states.
@@
Attractive polymer-arginine interactions approach an energetic minima at $\sim$0.5 nm (Fig. 1e), giving rise to the global minimum observed in the overall PMF. Polymer-Cl- interactions were observed on the order of thermal fluctuations, consistent with previous observations that Cl- ions are depleted from the local domain of hydrophobic solutes.104, 41
@@
Fig. 1h shows the change in each component upon unfolding in arginine solution relative to that observed in water. In Fig. 1h, the first $\Delta$ arises from the difference between folded and unfolded states (e.g., $\Delta E=\langle E_{u}\rangle-\langle E_{f}\rangle$), while the second $\Delta$ arises from the free energy difference between arginine solution ($\Delta E_{arg}$) and water ($\Delta E_{wat}$) (e.g., $\Delta\Delta E=\Delta E_{arg}-\Delta E_{wat}$). With increasing arginine concentration, we observed an increasingly favorable cavitation component for folding (Fig. 1h). Lin and Timasheff 105 previously connected protein stability with cavity formation and vapor-liquid surface tension. They proposed that the expansion of a protein-containing cavity requires more energy in the presence of an additive that increases surface tension at the cavity-solution interface. Experimentally, increasing arginine concentration has been observed to increase the vapor-liquid surface tension of aqueous solutions,106 which may explain the dependence of $\Delta\Delta G_{cav}$ observed in the present study.
@@
The polymer-arginine contribution favors the folded polymer state at 0.25 M and 0.5 M arginine concentrations relative to in pure water (Fig. 1h). However, at 1.0 M arginine concentrations, polymer-arginine interactions promote polymer unfolding. Together, these results indicate that neither direct nor indirect mechanistic hypotheses alone can describe the effects of arginine on hydrophobic polymer folding.
3.3 Mechanistic flip with increasing arginine concentration
To investigate potential competing effects of direct and indirect mechanisms, we combined the components of our PMF decomposition, delineating between those linked to direct effects (polymer-arginine and polymer-Cl-; $\Delta\Delta G_{dir}$) and indirect effects (cavitation and polymer-water; $\Delta\Delta G_{ind}$) of arginine (Fig. 2a).